40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7224     Entry time: Sat Aug 18 03:55:12 2012
Author: Yoichi 
Type: Summary 
Category: LSC 
Subject: X-arm locking again 

Tonight, I worked on the X-arm locking again. I did not have any significant progress, but observed several issues and will give some suggestions for future work here.

What I did tonight was basically re-alignment of the X-arm (because Rana touched the PZT mirrors for the Y-arm alignment, the X-arm alignment was screwed up). Then I measured the open loop gain. Of course it was almost identical to the one posted in this entry. It reminded myself of how small the phase bubble is. This means we have to finely adjust the gain to set the UGF at the right frequency, i.e. 100Hz. So I decided to do the signal normalization using the TRX power. Using the MC path method described here,  the appropriate normalization coefficient was determined to be 1.6, when the XARM gain is set to 0.05. Using burtgooey, I updated the burt snapshot used by the X-arm restore script.

Now I observed the following things:

When the normalization is used, the lock itself is stable, but the lock acquisition takes loner (i.e. fails more often).

I don't know the exact reason, but here is my guess: Usually, the error signal is divided by the square root of the transmitted power to widen the linear range of the PDH error signal. However, what I'm doing here is dividing the error signal with the power itself, not the sqrt. This might distort the error signal in a not-friendly-for-lock way ? I don't know.

I checked the c1lsc FE code. There seems to be the sqrt(TRX) and sqrt(TRY) signals computed in the code. However, these are not used for the normalization. 

Now, there are two requirements. When dragging the mirrors into the resonance, we want to normalize the error signal with sqrt(TRX). When the mirrors reach the resonance, the gain of the loop must be normalized by TRX. How do we smoothly connect those two states ? Someone should spend some time on this. Maybe I will work on this in Japan.  

We really need a time delay in the filter trigger

The automatic filter trigger is awesome. However, the [0^2:5^2] filter, which is an integrator, takes time to switch on and off. Every time the cavity passes by a resonance, this filter gets turned on and off slowly, giving some large transients. This transient combined with the bad coil balance of ETMX sometimes made the optical lever of ETMX crazy. This can be avoided by turning on this filter a few seconds after the power reaches the threshold. As Rana suggested, we should be able to put an arbitrary time delay to the filter trigger.

Someone should balance the coils

The coil balance of ETMX is bad and causing the above mentioned problem. I tweaked the coil balance by injecting a sinusoidal signal (10Hz) into ETMX pos and trying to minimizing the spectral peak in the optical lever signals. Of course, this is a cheesy work. Someone should put more serious effort on this.

A civilized interferometer should have an auto-alignment capability

After my alignment work, the X-arm power got to about 0.7. (This is probably because the MC transmission power has been low for the past 5 hours or so (attachment 1)).

In anyway, after the cavity locked to the TEM00 mode, the alignment has to be automatically improved by dithering. It is anachronism to sit down and click on the MEDM screen until the power gets big enough.

 

 

Attachment 1: MC_Trans.png  13 kB  | Hide | Hide all
MC_Trans.png
ELOG V3.1.3-